# Galaxy mergers classification using CNNs trained on Sérsic models, residuals and raw images

### v2

@@ 3.3.3 Evaluation metrics
4 Results
We show the performance of all three networks using the evaluation metrics described in Sect. 3. Evaluation metrics scores depend on the choice of a decision threshold as it changes the required score to classify galaxy as a merger. Here, we set the threshold for each CNN independently, but the change in metrics as a function of threshold was also examined in Fig. 6. For all CNNs, we choose the decision threshold at which precision equals recall to achieve a balanced trade-off between the rate of correctly identified positives and the rate of false positives, as both false positives and false negatives are equally undesirable. For CNN1 (raw images), CNN2 (model images), and CNN3 (residual images) it is 0.39, 0.47, and 0.49 respectively. CNN1 and CNN2 achieve their peak accuracy for these thresholds. The decline in accuracy after the peak is slower for CNN3 than CNN1 and CNN2. The overall performance across the entire redshift range for all networks is summarized in Table 2. The results show that the CNN1 performs the best in terms of accuracy - 74%. CNN2 and CNN3 have similar overall accuracy - correctly classifying about, respectively, 70% and 68% of all images. ROCs of all CNNs are shown in Fig. 8, and confusion matrices in Fig. 8. The overall best network is CNN1, followed by CNN2, and CNN3. CNN1 performs better in recognizing merging galaxies. 74.5% of them are correctly identified, while for non-merging galaxies, this number drops to 73.7%. CNN2 and CNN3 show similar performance. CNN2 correctly predicts 70.5% of mergers and 70.3% of non-mergers, while CNN3 correctly predicts 69.1% of mergers and 67.3% of non-mergers.
@@ 4.1 Redshift
4.2 Merger time
As seen by the higher fraction of correctly identified mergers, CNN1 achieves the best performance across most of the examined time intervals (-0.79 Gyr to 0.18 Gyr). CNN3 classifies mergers with similar precision as CNN1 in the later phase of the merging process (0.18 to 0.3 Gyr). As summarized in Table 3, CNN1 correctly classifies $79.3\pm 0.6\%$ of pre-mergers and $58.8\pm 1.5\%$ of post-mergers, CNN2 correctly classifies $76.6\pm 0.6\%$ of pre-mergers and $53.4\pm 1.6\%$ of post-mergers, and CNN3 correctly classifies $73.3\pm 0.6\%$ of pre-mergers and $55.5\pm 1.4\%$ of post-mergers. Reported values are the mean fractions of correctly classified post-mergers and pre-mergers for each model, with the error represented as the difference between the mean and the upper bound of the 95% confidence interval of 1000 bootstrap resamples.
@@ 4.3 Galaxy ellipticity
The metrics for each Sérsic ellipticity bin are shown in Fig. 13. We observe a high accuracy for low ellipticities across all networks, which then decreases and increases again around the bin (0.65, 0.7]. Precision decreases in the first three bins, then increases, reaching its maximum in the same bin. This results from the fact that the metrics are sensitive to class imbalance within a given bin. Higher ellipticities contain more mergers than non-mergers, so naturally, the number of TPs increases with the bin, while the number of FPs decreases. To mitigate the impact of this imbalance, the number of mergers and non-mergers was set equal in each bin. Each bin contains 500 mergers and 500 non-mergers. If the number of mergers or non-mergers in a bin is less than 500, the number of the less frequent type of galaxy is selected for that bin. Galaxies are randomly selected for each bin 20 times. Fig. 13 shows the average metrics along with their standard deviations. As a result, we now observe a significantly greater ability of all networks to recognize mergers at low ellipticities than at high ellipticities, as indicated by the higher precision and only slightly decreasing accuracy, which decreases significantly only for CNN3 at the highest ellipticities.
@@
Figure 13: Accuracy (top) and Precision (bottom) with standard deviation across Sérsic ellipticities for balanced bins. Each bin contains 500 mergers and 500 non-mergers. If the number of mergers or non-mergers in a bin is less than 500, the number of the less frequent type of galaxy is selected for that bin. Metrics for CNN1 are shown with a blue solid line and circles, for CNN2 with green dashed line and squares, for CNN3 with red dotted line and triangle facing upwards.
5 Discussion
5.1 CNNs comparison
Among the three networks, CNN1 performs best overall, achieving the highest classification accuracy of 74%. CNN1 has access to both residual and original image information, which appears to be most effective for distinguishing mergers from non-mergers. CNN2, which uses only the original images containing source position and shape information, performs slightly better (70% accuracy) than CNN3 (68% accuracy), which relies solely on residual features (see Table 2 for overall performance of all CNNs). This suggests that source position and shape information play a somewhat more significant role in classification than residual features alone. The absence of such information, especially features like double nuclei that may be removed through Sérsic profile subtraction, can confuse the network and reduce its ability to accurately distinguish between mergers and non-mergers.
@@ 5.1 CNNs comparison
Mergers with lower Sérsic profile ellipticity are more frequently correctly classified than those with higher ellipticity across all three CNNs, as shown in Fig. 13. This may be because merger-induced distortions are more prominent in galaxies with more symmetric profiles.
@@
5.2 Occlusion
Following Pearson et al. (2022), occlusion experiments were performed on five mergers and five non-mergers. This technique is used to visualize what parts of the image are important for classification, as CNN feature maps are very unintuitive in deeper layers. In Fig. 14 we choose mergers that were: a) classified as mergers by all three CNNs, b) classified as non-mergers by all three CNNs, c) classified as mergers by CNN1 and as non-mergers by CNN2 and CNN3, d) classified as mergers by CNN2, and as non-mergers by CNN1 and CNN3, and e) classified as mergers by CNN3 and as non-mergers by CNN1 and CNN2. We also take non-mergers in Fig. 15 that were: a) classified as non-mergers by all three CNNs, b) classified as mergers by all three CNNs, c) classified as non-mergers by CNN1 and as mergers by CNN2 and CNN3, d) classified as non-mergers by CNN2, and as mergers by CNN1 and CNN3, and e) classified as non-mergers by CNN3 and as mergers by CNN1 and CNN2. The occlusion experiment relies on covering the image with a box (we used 8×8 pixels), evaluating it, shifting the box one pixel to the right so it covers different pixels, evaluating it again, and repeating it all over until the box reaches the lower right corner of an image. The occluded images are treated as normal galaxy images by a network. For CNN1 and CNN3 we used a box filled with noise with the properties of the image background, and for CNN2 we used a box filled with 0 values. This is done to not introduce structures that were not previously seen by the network, as there are no empty chunks in original images and residual images and there are no noise structures on the model images. This technique allows us to create an importance heatmap, which represents the average classification when each pixel was occluded. For mergers, important parts of the image will have lower scores in the importance heatmap than less important parts, and for non-mergers, important parts of the image will have higher scores than less important parts.
@@ 5.2 Occlusion
It is important to acknowledge that these occlusion experiments were conducted on a relatively small sample due to computational constraints. As a result, while the observed patterns are compelling, they may not fully capture the diversity of behaviors across the entire dataset. Nevertheless, the consistency in CNN focus areas, particularly around primary galaxies and potential double nuclei, provides promising evidence that the networks are learning meaningful astrophysical features. We are therefore optimistic that this is indeed what the networks are attending to when making their classifications.
Figure 14: Occlusion experiment for mergers that were: a) classified as mergers by all three CNNs, b) classified as non-mergers by all three CNNs, c) classified as mergers by CNN1 and as non-mergers by CNN2 and CNN3, d) classified as mergers by CNN2, and as non-mergers by CNN1 and CNN3, and e) classified as mergers by CNN3 and as non-mergers by CNN1 and CNN2. First two left columns show CNN1 input and corresponding heatmap. Two middle columns show CNN2 input and corresponding heatmap. Last middle columns show CNN3 input and corresponding heatmap. Above each heatmap there is a prediction for the original input, before performing occlusion experminet. Thresholds for CNN1, CNN2, and CNN3 are 0.39, 0.47, and 0.49 respectively.
Figure 15: Occlusion experiment for non-mergers that were: a) classified as non-mergers by all three CNNs, b) classified as mergers by all three CNNs, c) classified as non-mergers by CNN1 and as mergers by CNN2 and CNN3, d) classified as non-mergers by CNN2, and as mergers by CNN1 and CNN3, and e) classified as non-mergers by CNN3 and as mergers by CNN1 and CNN2. First two left columns show CNN1 input and corresponding heatmap. Two middle columns show CNN2 input and corresponding heatmap. Last two columns show CNN3 input and corresponding heatmap. Above each heatmap there is a prediction for the original input, before performing occlusion experminet. Thresholds for CNN1, CNN2, and CNN3 are 0.39, 0.47, and 0.49 respectively.
@@ Acknowledgments
We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017417. W.J.P has been supported by the Polish National Science Center project UMO-2023/51/D/ST9/00147. A.P has been supported by the Polish National Science Center project UMO-2023/50/A/ST9/00579. L.E.S was supported by the Estonian Ministry of Education and Research (grant TK202), Estonian Research Council grant (PRG1006), and the European Union’s Horizon Europe research and innovation programme (EXCOSM, grant No. 101159513). B.M.B and L.W are supported by the project ‘Clash of the Titans: deciphering the enigmatic role of cosmic collisions’ (with project number VI.Vidi.193.113 of the research programme Vidi, which is (partly) financed by the Dutch Research Council (NWO).

### v3

@@ III.3.3 Evaluation metrics
IV Results
We show the performance of all three networks using the evaluation metrics described in Sect. III. Evaluation metrics scores depend on the choice of a decision threshold as it changes the required score to classify galaxy as a merger. Here, we set the threshold for each CNN independently, but the change in metrics as a function of threshold was also examined in Fig. 6. For all CNNs, we choose the decision threshold at which precision equals recall to achieve a balanced trade-off between the rate of correctly identified positives and the rate of false positives, as both false positives and false negatives are equally undesirable. For CNN1 (raw images), CNN2 (model images), and CNN3 (residual images) it is 0.39, 0.47, and 0.49 respectively. CNN1 and CNN2 achieve their peak accuracy for these thresholds. The decline in accuracy after the peak is slower for CNN3 than CNN1 and CNN2. The overall performance across the entire redshift range for all networks is summarized in Table 2. The results show that the CNN1 performs the best in terms of accuracy - 74%. CNN2 and CNN3 have similar overall accuracy - correctly classifying about, respectively, 70% and 68% of all images. ROCs of all CNNs are shown in Fig. 8, and confusion matrices in Fig. 8. The overall best network is CNN1, followed by CNN2, and CNN3. CNN1 performs better in recognizing merging galaxies. 74.5% of them are correctly identified, while for non-merging galaxies, this number drops to 73.7%. CNN2 and CNN3 show similar performance. CNN2 correctly predicts 70.5% of mergers and 70.3% of non-mergers, while CNN3 correctly predicts 69.1% of mergers and 67.3% of non-mergers.
@@ IV.1 Redshift
IV.2 Merger time
As seen by the higher fraction of correctly identified mergers, CNN1 achieves the best performance across most of the examined time intervals (-0.79 Gyr to 0.18 Gyr). CNN3 classifies mergers with similar precision as CNN1 in the later phase of the merging process (0.18 to 0.3 Gyr). As summarized in Table 3, CNN1 correctly classifies $79.3\pm 0.6\%$ of pre-mergers and $58.8\pm 1.5\%$ of post-mergers, CNN2 correctly classifies $76.6\pm 0.6\%$ of pre-mergers and $53.4\pm 1.6\%$ of post-mergers, and CNN3 correctly classifies $73.3\pm 0.6\%$ of pre-mergers and $55.5\pm 1.4\%$ of post-mergers. Reported values are the mean fractions of correctly classified post-mergers and pre-mergers for each model, with the error represented as the difference between the mean and the upper bound of the 95% confidence interval of 1000 bootstrap resamples. The substantially lower fractions of correctly classified post‑mergers compared to pre‑mergers for all three CNNs are likely influenced by the relative scarcity of post‑mergers in the training data. As seen in Table 1, the underlying mock sample contains fewer post‑mergers than pre‑mergers, which tends to bias the networks toward the better‑represented pre‑merger morphologies. This imbalance, in combination with the intrinsically fainter features expected in post‑mergers, probably contributes to the systematically poorer performance of all models in this regime.
@@ IV.3 Galaxy ellipticity
The metrics as a function of Sérsic ellipticity, shown in Fig. 12, exhibit consistently high accuracy at low ellipticities for all three networks, followed by a general decline toward intermediate ellipticities and a mild recovery at (0.45, 0.65] bins. Precision rises more steadily: after an initial drop in the (0, 0.15] bins, it increases across most of the ellipticity range, reaching its peak around the (0.65, 0.70] bin before decreasing toward the highest ellipticities. These trends reflect the sensitivity of both metrics to the varying merger fraction within each bin - higher ellipticity bins contain relatively more mergers, which leads to more true positives and fewer false positives. These trends suggest that ellipticity provides useful discriminatory power for merger classification.
@@
V Discussion
V.1 CNNs comparison
Among the three networks, CNN1 performs best overall, achieving the highest classification accuracy of 74%. CNN1 has access to both shape and position information, as well as any residual features present in the data, which appears to be most effective for distinguishing mergers from non-mergers. CNN2, which uses only source position and shape information, performs slightly better (70% accuracy) than CNN3 (68% accuracy), which relies solely on residual features (see Table 2 for overall performance of all CNNs). This suggests that source position and shape information play a somewhat more significant role in classification than residual features alone. The absence of such information, especially features like double nuclei that may be removed through Sérsic profile subtraction, can confuse the network and reduce its ability to accurately distinguish between mergers and non-mergers.
@@ V.1 CNNs comparison
V.2 Occlusion
Following Pearson et al. (2022), occlusion experiments were performed on five mergers and five non-mergers from the test set and the $0.76<z<1$ redshift range. This technique is used to visualize what parts of the image are important for classification, as CNN feature maps are very unintuitive in deeper layers. The occlusion experiment relies on covering the image with a box (we used 8×8 pixels), evaluating the models’ predicted class probability after occluding, shifting the box one pixel to the right so it covers different pixels, evaluating the class again, and repeating it all over until the box reaches the lower right corner of an image. The occluded images are treated as normal galaxy images by a network. For CNN1 and CNN3 we used a box filled with noise with the properties of the image background, and for CNN2 we used a box filled with 0 values. This is done to not introduce structures that were not previously seen by the network, as there are no empty chunks in original images and residual images and there are no noise structures on the model images. This technique allows us to create an importance heatmap, which represents the average classification when each pixel was occluded. For mergers, important parts of the image will have lower scores in the importance heatmap than less important parts, and for non-mergers, important parts of the image will have higher scores than less important parts. In Fig. 13 we choose mergers that were: a) classified as mergers by all three CNNs, b) classified as non-mergers by all three CNNs, c) classified as mergers by CNN1 and as non-mergers by CNN2 and CNN3, d) classified as mergers by CNN2, and as non-mergers by CNN1 and CNN3, and e) classified as mergers by CNN3 and as non-mergers by CNN1 and CNN2. We also take non-mergers in Fig. 14 that were: a) classified as non-mergers by all three CNNs, b) classified as mergers by all three CNNs, c) classified as non-mergers by CNN1 and as mergers by CNN2 and CNN3, d) classified as non-mergers by CNN2, and as mergers by CNN1 and CNN3, and e) classified as non-mergers by CNN3 and as mergers by CNN1 and CNN2.
@@ V.2 Occlusion
It is important to acknowledge that these occlusion experiments were conducted on a relatively small sample due to computational constraints. As a result, while the observed patterns are compelling, they may not fully capture the diversity of behaviours across the entire dataset. Nevertheless, the consistency in CNN focus areas, particularly around primary galaxies and potential double nuclei, provides promising evidence that the networks are learning meaningful astrophysical features. We are therefore optimistic that this is indeed what the networks are attending to when making their classifications.
Our occlusion results show strong qualitative consistency with the image-only network in Pearson et al. (2022). In both cases the primary galaxy is the dominant feature driving the classification, with the lowest occlusion scores typically occurring where the central light concentration or potential double nuclei reside. Similar to Pearson et al. (2022), our networks also show that bright secondary galaxies can influence the classification, sometimes being interpreted as potential companions even when they are unrelated. A notable difference is that our CNNs, especially the residual-based one, show a clearer sensitivity to faint structures around the primary galaxy, while Pearson et al. (2022) reported weaker evidence for such features in their occlusion maps. Overall, however, the importance patterns are broadly aligned, indicating that both sets of networks rely on comparable visual cues when identifying mergers.
Generating an occlusion map took 3 minutes per image on an RTX 4090.
Figure 13: Occlusion experiment for mergers that were: a) classified as mergers by all three CNNs, b) classified as non-mergers by all three CNNs, c) classified as mergers by CNN1 and as non-mergers by CNN2 and CNN3, d) classified as mergers by CNN2, and as non-mergers by CNN1 and CNN3, and e) classified as mergers by CNN3 and as non-mergers by CNN1 and CNN2. First two left columns show CNN1 input and corresponding heatmap. Two middle columns show CNN2 input and corresponding heatmap. Last middle columns show CNN3 input and corresponding heatmap. Above each heatmap there is a prediction for the original input, before performing occlusion experiment. Thresholds for CNN1, CNN2, and CNN3 are 0.39, 0.47, and 0.49 respectively. Pixels where occlusion decreases the predicted merger probability are shown in darker red and correspond to regions of high feature importance, whereas pixels where occlusion increases the merger probability appear in lighter colors, indicating low importance for the classification
Figure 14: Occlusion experiment for non-mergers that were: a) classified as non-mergers by all three CNNs, b) classified as mergers by all three CNNs, c) classified as non-mergers by CNN1 and as mergers by CNN2 and CNN3, d) classified as non-mergers by CNN2, and as mergers by CNN1 and CNN3, and e) classified as non-mergers by CNN3 and as mergers by CNN1 and CNN2. First two left columns show CNN1 input and corresponding heatmap. Two middle columns show CNN2 input and corresponding heatmap. Last two columns show CNN3 input and corresponding heatmap. Above each heatmap there is a prediction for the original input, before performing occlusion experminet. Thresholds for CNN1, CNN2, and CNN3 are 0.39, 0.47, and 0.49 respectively. Pixels where occlusion increases the predicted non-merger probability are shown in darker red and correspond to regions of high feature importance, whereas pixels where occlusion decreases the merger probability appear in lighter colors, indicating low importance for the classification.
@@ Acknowledgments
We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017417. We also thank Tomasz Łaguz for providing access to an RTX 4090, which enabled on-demand training, testing, and evaluation. W.J.P has been supported by the Polish National Science Center project UMO-2023/51/D/ST9/00147. A.P has been supported by the Polish National Science Center project UMO-2023/50/A/ST9/00579. L.E.S was supported by the Estonian Ministry of Education and Research (grant TK202), Estonian Research Council grant (PRG1006), and the European Union’s Horizon Europe research and innovation programme (EXCOSM, grant No. 101159513). B.M.B and L.W are supported by the project ‘Clash of the Titans: deciphering the enigmatic role of cosmic collisions’ (with project number VI.Vidi.193.113 of the research programme Vidi, which is (partly) financed by the Dutch Research Council (NWO).
Table 4: The CNN architecture.
Normalized input images (raw, model, residual), range [0,1]
Conv2D: 128 filters, 13×13 kernel
ReLU, Batch Normalization, Dropout (20%), MaxPooling 2×2
Conv2D: 128 filters, 11×11 kernel
ReLU, Batch Normalization, Dropout (20%), MaxPooling 2×2
Conv2D: 128 filters, 9×9 kernel
ReLU, Batch Normalization, Dropout (20%), MaxPooling 2×2
Flatten
FC1: 512 neurons
ReLU, Batch Normalization, Dropout (20%)
FC2: 128 neurons
ReLU, Batch Normalization, Dropout (20%)
Output layer: 1 neuron
Sigmoid
Loss: Binary cross-entropy
Optimizer: Adam ($\alpha=3\cdot 10^{-4}$)
Epochs: 50
8 966 273 trainable parameters